Use any method to solve the problem: 4y - y = ex/2 + 4,
Answer #1

to solve the thod problem me use any uy - y +4. sol H2 4- y е +4. 0 y=D²y, yady y and yl values substistute inequation 21y(x) = geta x + ez Yele) = Gette emne particular integral (p. I). Ypres) x12 +4: (4D2-1) p = 지오 + Ypa 14.) 4D²-1 402-1 ( anni216 xe Yp (x)= -4. 4. ५ particular solution pro particular integral is a ez Oprava 4 4 The general Solution is sum of complem